OpenStudy (anonymous):

i'll do anything for the answer :) simplify the radical expression sqrt 3+4 sqrt 3

OpenStudy (btaylor):

Think of it as \(1 \cdot \sqrt{3} + 4 \cdot \sqrt{3}\). What could you factor out of both terms?

OpenStudy (anonymous):

Thank you <3 :)

OpenStudy (btaylor):

@llhome please don't just give answers. It ruins OS for those of us who actually try to follow the Code of Conduct.